# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

### v3

@@ Discussion.
In contrast, commercial models like GPT-4o Mini and Claude 3.5 achieved high performance across all settings, with only marginal benefits from added retrieval. This behavior is likely due to their extensive pretraining on diverse codebases, which may already include material similar to what our dataset provides. Consequently, the impact of external retrieval is diminished and can even introduce noise if not precisely targeted. These findings emphasize the importance of retrieval quality over quantity in RAG-based code generation. The consistent advantage of partial over full context suggests that carefully curated, relevant examples are more beneficial than exhaustive inclusion.
@@ VI Conclusion
In this paper, we introduce PennyLang, a high-quality instruction–response dataset tailored for PennyLane-based quantum code generation. We present our data collection and prepossessing methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples. To evaluate its practical utility, we develop a RAG evaluation framework and benchmark four state-of-the-art LLMs (GPT-4o Mini, Claude 3.5 Sonnet, Qwen2.5-7B, and LLaMa 4-17B Maverick) under both RAG and non-RAG settings. Our findings show that PennyLang, when paired with RAG, substantially boosts performance for open-source models, while proprietary models exhibit minimal or negative gains with naive full-context retrieval, likely because their pre-training already covers overlapping PennyLane-related knowledge. This further highlights the importance of calibrated retrieval budgets, as moderate settings (e.g., 75%) often outperform full context by balancing signal and noise. We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development. Future directions include training lightweight instruction-tuned models on this dataset, optimizing retrieval strategies across different model types, and extending to cross-framework generalization between PennyLane, Qiskit, and Cirq.

### v4

@@ Discussion.
In contrast, commercial models like GPT-4o Mini, Claude 3.5 Sonnet, Gemini 2.5 Flash, Claude Haiku, and GPT-5 Mini achieved high performance across all settings, with only marginal benefits from added retrieval. This behavior is likely due to their extensive pretraining on diverse codebases, which may already include material similar to what our dataset provides. Consequently, the impact of external retrieval is diminished and can even introduce noise if not precisely targeted. These findings emphasize the importance of retrieval quality over quantity in RAG-based code generation. The consistent advantage of partial over full context suggests that carefully curated, relevant examples are more beneficial than exhaustive inclusion.
@@ VI Conclusion
In this paper, we introduce PennyLang, a high-quality instruction–response dataset tailored for PennyLane-based quantum code generation. We present our data collection and preprocessing methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples. To evaluate its practical utility, we develop a RAG evaluation framework and benchmark seven state-of-the-art LLMs (both open-source and commercial models) under both RAG and non-RAG settings. Our findings show that PennyLang, when paired with RAG, substantially boosts performance for open-source models, while proprietary models exhibit minimal or negative gains with naive full-context retrieval, likely because their pre-training already covers overlapping PennyLane-related knowledge. This further highlights the importance of calibrated retrieval budgets, as moderate settings (e.g., 75%) often outperform full context by balancing signal and noise. We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development. Future directions include training lightweight instruction-tuned models on this dataset, improving the filtering pipeline for continuously collecting high-quality new code snippets into the database, optimizing retrieval strategies across different model types, and extending to cross-framework generalization between PennyLane, Qiskit, and Cirq.
Data Availability
The dataset proposed in this paper is publicly available on GitHub at https://github.com/eBrain4Everyone/Pennylang.